# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v5

@@ 2.2.2 RED prior
    Jacobian symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{n}$.
@@ 2.2.3 Implicit Gradient RED
The nonexpansive assumption has theoretically played a crucial role in PnP prior and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49]:
@@
By assumption conditions, the RED term $g_{\text{red}}$ is convex, and $\nabla g_{\text{red}}=I-D_{\sigma}=R$. Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser. Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex. Applying Proposition 2.5, together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=\mathbf{x}$, there exists a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required. From the perspective of monotone operator theory, Lemma 2.8 [2, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
@@
Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$. Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator. Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudocontractive denoisers [57], have made the nonexpansive assumption more realistic in practice. Constructing the FNE denoiser $D_{\sigma}$ is feasible, similarly, the residual $R=I-D_{\sigma}$ can be FNE in (7) with $\mathcal{M}=I$. Both can be proximal operators of implicit regularizers.
@@ 3 Halpern-type preconditioned proximal point (HPPP)
Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$ [26] is an implicitly regularized method for finding a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$. When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$. Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
@@ 3.1 Convergence analysis
(12) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
@@
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfies conditions (i)-(iii) of Theorem 3.1. Compared with [3], we have provided an alternative method to obtain the limit point as the $\mathcal{M}$-projection of the initial point. The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$. By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. According to Theorem 3.1, we obtain the following proposition.
@@
Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], and hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$. Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. This completes the proof of Proposition 3.3.
@@
We compare Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same, except for the additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption, especially in applications to splitting algorithms, and is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly [31], where $U$ is a closed convex cone and $V$ is a closed hyperplane. Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence. Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
@@
As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
@@
and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$. By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$. Indeed, if by contraction, $\mathbf{u}^{k}\to\mathbf{0}$, then
@@
Regarding the choice of the anchor point, we can use the degraded observed image, the denoised image, or the restarted technique (see Algorithm 2). The restarted technique is an effective method in machine learning [41] and LP [17, 38].
@@ 3.2 Convergence rate
Based on the following technical lemma, Sabach and Shtern [50] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
Lemma 3.9 ([50]).
@@
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3]. By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE. Let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, and then HPPP is equivalent to the following reduced algorithm:
@@
By [36, Theorem 2.1] and [52, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
@@
    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$,
@@ 3.3 GraRED-HP3
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points. Under the ideal conditions of RED and with assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$; then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], and we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17). Set $\mathbf{K}=I$; the implicit gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3. Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers. According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
@@ 3.3.1 Acceleration for PnP methods
PnP-ADMM (see Algorithm 1) is a well-known method for solving (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$. When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration. Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9]. The related GraRED-P3 algorithm is shown in Algorithm 4. Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
@@
In [28], a worst-case convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{k}}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM. In contrast, GraRED-HP3 obtains an accelerated convergence rate of $\mathcal{O}\left(\frac{1}{k}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.

### v6

@@ 2.2.2 RED prior
    Jacobian symmetry [REDClarifications2019]: $\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{n}$.
@@ 2.2.3 Implicit Gradient RED
The nonexpansive assumption has theoretically played a crucial role in PnP prior and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [ryu2019plug]:
@@
By assumption conditions, the RED term $g_{\text{red}}$ is convex, and $\nabla g_{\text{red}}=I-D_{\sigma}=R$. Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser. Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex. Applying Proposition 2.5, together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [beck2017first, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=\mathbf{x}$, there exists a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required. From the perspective of monotone operator theory, Lemma 2.8 [BCombettes, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
@@
Moreover, by [BCombettes, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$. Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator. Recent developments, such as spectral normalization [ryu2019plug, pmlr-v202-delattre23a], learning maximally monotone operators [MMO2021], learning FNE denoisers [building_FNE, bredies2024learningfirmlynonexpansiveoperators], and learning pseudocontractive denoisers [Wei2024], have made the nonexpansive assumption more realistic in practice. Constructing the FNE denoiser $D_{\sigma}$ is feasible, similarly, the residual $R=I-D_{\sigma}$ can be FNE in (7) with $\mathcal{M}=I$. Both can be proximal operators of implicit regularizers.
@@ 3 Halpern-type preconditioned proximal point (HPPP)
Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [he2024convergence]. The Halpern iteration $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$ [halpern1967fixed] is an implicitly regularized method for finding a particular fixed point [Diakonikolas2020], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [halpern1967fixed, lions1977approximation, Wittmann1992, 2002Iterative, xu2004viscosity, Qi2021, he2024convergence], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$. When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$. Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
@@ 3.1 Convergence analysis
[…]
@@
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfies conditions (i)-(iii) of Theorem 3.1. Compared with [Bauschke2023], we have provided an alternative method to obtain the limit point as the $\mathcal{M}$-projection of the initial point. The work [Bauschke2023] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$. By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. According to Theorem 3.1, we obtain the following proposition.
@@
Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [BCombettes, Proposition 20.38], and hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$. Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. This completes the proof of Proposition 3.3.
@@
We compare Theorem 3.1, Proposition 3.3 with [Bredies2022, Theorem 2.9] and [Bredies2022, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same, except for the additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption, especially in applications to splitting algorithms, and is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly [HUNDAL200435], where $U$ is a closed convex cone and $V$ is a closed hyperplane. Building on Hundal’s counterexample, [DR_weak_convergence2020] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence. Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
@@
As proved in [DR_weak_convergence2020], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
@@
and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$. By [Bredies2022, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$. Indeed, if by contraction, $\mathbf{u}^{k}\to\mathbf{0}$, then
@@
Regarding the choice of the anchor point, we can use the degraded observed image, the denoised image, or the restarted technique (see Algorithm 2). The restarted technique is an effective method in machine learning [Park2022] and LP [chen2024hprlpimplementationhprmethod, lu2024restartedhalpernpdhglinear].
@@ 3.2 Convergence rate
Based on the following technical lemma, Sabach and Shtern [Sabach2017] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
Lemma 3.9 ([Sabach2017]).
@@
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [Bredies2022, Proposition 2.3]. By [Bredies2022, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE. Let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, and then HPPP is equivalent to the following reduced algorithm:
@@
By [lieder2021convergence, Theorem 2.1] and [Sun2024, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
@@
[…]
@@ 3.3 GraRED-HP3
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points. Under the ideal conditions of RED and with assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$; then $\phi^{*}=sg^{*}$ [beck2017first, Theorem 4.14], and we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17). Set $\mathbf{K}=I$; the implicit gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3. Even if the conditions of RED are difficult to meet [REDClarifications2019], recent developments [ryu2019plug, building_FNE, MMO2021, pmlr-v202-delattre23a, BUI2020124315, bredies2024learningfirmlynonexpansiveoperators, Wei2024] can learn FNE denoisers. According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
@@ 3.3.1 Acceleration for PnP methods
PnP-ADMM (see Algorithm 1) is a well-known method for solving (1), which can be written into an equivalent DRS form [ryu2019plug, hurault2022proximal], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$. When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration. Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [Bredies2015, Bredies2022]. The related GraRED-P3 algorithm is shown in Algorithm 4. Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
@@
In [He2015], a worst-case convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{k}}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM. In contrast, GraRED-HP3 obtains an accelerated convergence rate of $\mathcal{O}\left(\frac{1}{k}\right)$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.